# Leveraging Sidewalk Robots for Walkability-Related Analyses

### v2

@@ Abstract
Walkability is a key component of sustainable urban development, while collecting detailed data on sidewalks (or pedestrian infrastructures) remains challenging due to the high costs and limited scalability of traditional methods. Sidewalk delivery robots, increasingly deployed in urban environments, offer a promising solution to these limitations. This paper explores how these robots can serve as mobile data collection platforms, capturing sidewalk-level features related to walkability in a scalable, automated, and real-time manner. A sensor-equipped robot was deployed on a sidewalk network at KTH in Stockholm, completing 101 trips covering 900 segment records. From the collected data, different typologies of features are derived, including robot trip characteristics (e.g., speed, duration), sidewalk conditions (e.g., width, surface unevenness), and sidewalk utilization (e.g., pedestrian density). Their walkability-related implications were investigated with a series of analyses. The results demonstrate that pedestrian movement patterns are strongly influenced by sidewalk characteristics, with higher density, reduced width, and surface irregularity associated with slower and more variable trajectories. Notably, robot speed closely mirrors pedestrian behavior, highlighting its potential as a proxy for assessing pedestrian dynamics. The proposed framework enables continuous monitoring of sidewalk conditions and pedestrian behavior, contributing to the development of more walkable, inclusive, and responsive urban environments.
@@ 1 Introduction
Whether operational or perceived, walkability is increasingly recognized as a vital aspect of sustainable urban development. To develop more walkable cities, one must first unravel the complex pedestrian mobility patterns and interactions with existing infrastructure (Rhoads et al.,, 2023). Thus, increasing efforts are being made in assessing the usage of pedestrian infrastructure, identifying critical features (e.g., pedestrian density, pedestrian speed), and understanding the complex pedestrian mobility patterns and interactions (Fruin,, 1970; Maghelal and Capp,, 2011; Dragović et al.,, 2023). Currently, pedestrian flow data is typically collected through infrared (IR) break-beam sensors and cameras placed at specific locations across the city (Moussaïd et al.,, 2010; Alahi et al.,, 2014; Carter et al.,, 2020). While this type of solution can offer granular data, it is typically difficult to adopt it on a broad scale due to its high costs and privacy issues (e.g., General Data Protection Regulation in Europe). More recently, street-level imagery (e.g., Google Street View) has been explored as an alternative source, where pedestrians can be detected and counted with computer vision techniques (Liu et al.,, 2023; Chen et al.,, 2020). These imagery-based approaches are less costly to deploy and scalable across large areas, but they suffer from limited temporal resolution, since imagery is updated infrequently and cannot provide continuous measurements. Other studies focusing on the larger-scale analysis of urban mobility patterns (e.g., origins and destinations, mode choice, trip features) rely on GPS, wireless and cellular networks, and social media geotag records (Leslie et al.,, 2007; Kung et al.,, 2014; D’Orso and Migliore,, 2020; Gao et al.,, 2014). These solutions offer a good overview of pedestrian mobility at an aggregate level, but they cannot capture more complex phenomena such as, pedestrian flows, trajectories, and interactions with other road users. Surveys and audits are also used for collecting qualitative data on perceived walkability (Steve and Shane,, 2011; Cambra and Moura,, 2020; Chen et al.,, 2024) and quantitative measures of the built environment to develop walkability indicators (Bartzokas-Tsiompras et al.,, 2021; Lee et al.,, 2020). Although these methods provide useful insights, they are time-consuming, resource-intensive, and difficult to scale, posing a major bottleneck to advancing the knowledge of walkability in research.
@@
This study does not aim to define walkability. Instead, it demonstrates how sidewalk robots can be leveraged to automatically capture measurable walkability-related features and use them to assess pedestrian mobility in practice. The contributions of this study are threefold. First, it introduces robots as a novel data collection instrument, capable of overcoming limitations of traditional methods which prevent their application and scalability. As sidewalk delivery robots become more common in urban environments, their routine movements offer a practical and cost-efficient means to evaluate pedestrian mobility real-time. Second, by grouping different operational, walkability-related factors into robot trip features, sidewalk condition features, and sidewalk utilization features, the study establishes a systematic framework for extracting and analyzing data. Third, it provides empirical evidence that robot-based measurements can both confirm established relationships in pedestrian dynamics and reveal context-specific patterns, such as how width, density, and surface quality shape walking speed and movement styles. In the long run, this robot-based method can help build richer and more inclusive datasets that complement traditional audits and fixed sensors.
@@ 2 Related Work
While these studies offer a broad conceptual and empirical understanding of walkability, this study focuses on the operational features, specific environmental factors that directly affect the functional experience of walking. Rather than a systematic survey of the entire walkability literature, we summarize relevant across complementary fields, including urban planning, transport geography, pedestrian mobility studies, and traffic flow theory. The selected studies explicitly examined the relation between the built environment and pedestrian movements, provided measurable indicators at the sidewalk level, and offered empirical insights. On this basis , we identified twelve relevant factors from an operational perspective: 1) static obstacles; 2) dynamic obstacles; 3) walkway width; 4) walkway quality; 5) sidewalk slope; 6) crosswalk distance; 7) crosswalk type; 8) potential for conflicts; 9) vehicle speed; 10) traffic volume; 11) driveways; 12) supporting facilities. Among these, factors 6 to 12 are mainly related to safety and crossing conditions. Each feature may correspond to multiple indicators or measurements across different studies.
@@ 3.1 Experimental setup
From September 18, 2024, to March 24, 2025, the robot completed 101 trips, covering 900 records of sidewalk segments in total. Trips were conducted on different days and at various times between 08:00 and 18:00, with particular focus on peak hours. During each trip, onboard sensors continuously recorded raw data, including GNSS points, robot velocity and heading, IMU readings, sidewalk width, ambient brightness, and anonymized pedestrian data. The experiments were conducting during months of higher pedestrian activity on campus, ensuring the dataset’s relevance. This dataset was processed to derive three categories of walkability-related features: robot trip features, sidewalk condition features, and sidewalk utilization features. These extracted features offer valuable insights into the quality of walking infrastructure and spatio-temporal pedestrian patterns. An overview of these features is presented in Table 2, while an interactive dashboard [^1] https://ismir-sidewalk-mobility-data-dashboard.streamlit.app/ visualizes sidewalk usage and characteristics for the study area.
@@ 3.4 Sidewalk utilization features
4 Feature-Based Analysis
We perform a series of analyses on the extracted sidewalk features to investigate their walkability-related implications. Subsection 4.1 presents an exploratory analysis of sidewalk condition features, highlighting variation in surface quality, width, and slope across different segments. Subsection 4.2 shifts the focus to aggregated pedestrian utilization patterns like speed and density relations. Building on these analyses, Subsection 4.3.1 applies clustering to identify distinct pedestrian behavior categories and analyze their relation to environmental and infrastructure characteristics. Finally, Subsection 4.3.2 employs regression to investigate the potential influence of sidewalk operational features on pedestrian movement.
@@ 4.1 Sidewalk Condition Variability Across the Network
In contrast, the irregularity index exhibits higher variability across the network, with Segment 7 displaying a significant spike due to frequent paving pattern changes (see Figure 7(b)). When the robot travels through this segment, it encounters four abrupt paving transitions within a short distance, including a cycling lane in between. Despite Segment 7 recording the highest irregularity, its unevenness is moderate, demonstrating that the proposed IMU-based method can effectively discriminate between these two characteristics. Segment 8 exhibits the second-highest irregularity score, reflecting the presence of a manhole cover and a storm drain in that location. On the other hand, Segments 2 and 3 show minimal irregularity values, suggesting smoother and more regular surfaces. Together, these results confirm that our proposed unevenness and irregularity metrics accurately describe actual sidewalk conditions, providing valuable insights for assessing pedestrian comfort and sidewalk robot mobility.
The distributions of detected average effective walkway width during the robot’s trips across various sidewalks are shown in Figure 7(c). Segments 1, 2, 3, 4, and 7 have highly concentrated width distributions, while the remaining segments display a wider spread in values. This variation can be largely attributed to differences in detection accuracy, which depends on both the physical characteristics of the sidewalk and the robot’s sensor limitations.
Segments 1, 2, 3, 4, and 7 are relatively narrow, with a width of under 3 meters. This makes boundary detection more consistent and reliable. In contrast, the other segments are situated in more open spaces rather than traditional sidewalk corridors, with widths ranging from approximately 5.5 to 10 meters. These areas often lack clearly defined edges, leading to increased variability in width detection.
The robot’s sensor limitation refers to the mounted stereo camera. Although it offers a 120-degree field of view, in practice it captures sidewalk information up to a maximum of about 7 meters ahead when centered. Because the robot frequently steers to avoid pedestrians or obstacles, the effective width detected may be underestimated and fluctuate in wider areas. These results suggest that the robot’s width measurements are more reliable in narrow, well-defined sidewalk segments, while detection accuracy decreases in broader or less structured environments. Additionally, some variation detected within the narrow, well-defined segments may result from temporary obstacles present during the robot’s passage.
Figure 7(d) presents the absolute slope values for each sidewalk segment, representing the steepness of each segment regardless of uphill or downhill direction. The highest absolute slope is observed in Segment 7, while Segments 1 and 2 exhibit the lowest slope values. This measurement can well reflect the varying elevation of different sidewalks, which is important for assessing walkability and planning sidewalk robot navigation.
@@ 4.2 Sidewalk Utilization and Density-Speed Dynamics
Average pedestrian speed and pedestrian density are estimated during each robot traversal of the sidewalk segments. The distributions of these two features across different sidewalks are presented in Figure 8(a) and Figure 8(b). The results reveal notable variability in pedestrian behavior across segments. Segments 3 and 4 show broader speed distributions, indicating a wider range of pedestrian movement speeds, which may be influenced by mixed usage or variable flow conditions. In contrast, Segments 6 and 9 exhibit more concentrated speed distributions, suggesting more uniform pedestrian movement in those areas.
Pedestrian density also varies significantly across sidewalks. For instance, although segments 6 and 7 are located on opposite sides of the same road, segment 6 exhibited the highest density levels, whereas segment 7 maintained substantially lower densities. As shown in Figure 7(c), their average sidewalk widths are approximately 6 m and 2 m, respectively, underscoring the pronounced difference in pedestrian volumes. These patterns align with on-site observations (Figure 3(a)): segment 6 serves as the main access corridor connecting the metro and light rail station exits to the campus entrance, and it is lined with convenience stores and restaurants. By contrast, segment 7 lacks notable points of interest, resulting in limited pedestrian activity.
@@
Compared to prior studies on pedestrian flow, the observed pedestrian density in our dataset is significantly lower, with a maximum of approximately 0.34 pedestrians per square meter, whereas previous empirical studies report densities reaching up to 10 pedestrians per square meter. Conversely, the maximum observed pedestrian speed in our data reaches around 2.5 m/s, exceeding the typical upper bound of around 1.7 m/s found in earlier work. These differences can be attributable to the characteristics of the study area: an open outdoor sidewalk network at a university campus, where most pedestrians are students or staff engaged in purposeful movement (e.g., commuting between classes). Measurement errors due to the robot’s onboard hardware and perception-limited capabilities may also contribute to such discrepancy. The robot relies on real-time computing pedestrian detection algorithms (Jocher et al.,, 2023), which are constrained by computational power and may not achieve the same level of precision as offline or stationary sensing systems.
The overall trend of Figure 9(a) aligns with established pedestrian flow theory, exhibiting an inverse relationship whereby increasing pedestrian density is associated with decreasing walking speed. At low densities (below 0.1 peds/m2), pedestrian speeds vary widely from below 0.5 m/s to over 2 m/s, indicating heterogeneous pedestrian behavior influenced by real-world factors such as individual urgency and route preference. This contrasts with the more constrained speed range (around 1 m/s to under 2 m/s) reported in Figure 9(b) for similar low-density conditions. As density increases beyond 0.15 peds/m2, pedestrian speeds tend to converge below 1.5 m/s, suggesting the onset of congestion and greater interaction among pedestrians. Furthermore, there are notable differences among sidewalk segments. For instance, Segment 4 exhibits a broader spread of speeds across multiple density levels, while Segment 1 consistently supports higher walking speeds at moderate densities. These observations highlight how pedestrian behavior varies across different urban segments. Overall, the results support the applicability of the fundamental diagram to outdoor urban sidewalk environments and underscore the influence of segment-level characteristics on pedestrian flow dynamics.
@@ 6 Conclusion
Acknowledgments
This study was funded by Digital Futures under the project: ”Investigating Sidewalks’ Mobility and Improving it with Robots (ISMIR).”

### v3

@@ Abstract
Walkability is a key component of sustainable urban development. In walkability studies, collecting detailed pedestrian infrastructure data remains challenging due to the high costs and limited scalability of traditional methods. Sidewalk delivery robots, increasingly deployed in urban environments, offer a promising solution to these limitations. This paper explores how these robots can serve as mobile data collection platforms, capturing sidewalk-level features related to walkability in a scalable, automated, and real-time manner. A sensor-equipped robot was deployed on a sidewalk network at KTH in Stockholm, completing 101 trips covering 900 segment records. From the collected data, different typologies of features are derived, including robot trip characteristics (e.g., speed, duration), sidewalk conditions (e.g., width, surface unevenness), and sidewalk utilization (e.g., pedestrian density). Their walkability-related implications were investigated with a series of analyses. The results demonstrate that pedestrian movement patterns are strongly influenced by sidewalk characteristics, with higher density, reduced width, and surface irregularity associated with slower and more variable trajectories. Notably, robot speed closely mirrors pedestrian behavior, highlighting its potential as a proxy for assessing pedestrian dynamics. The proposed framework enables continuous monitoring of sidewalk conditions and pedestrian behavior, contributing to the development of more walkable, inclusive, and responsive urban environments.
@@ 1 Introduction
Whether operational or perceived, walkability is increasingly recognized as a vital aspect of sustainable urban development. To develop more walkable cities, one must first unravel the complex pedestrian mobility patterns and interactions with existing infrastructure (Rhoads et al.,, 2023). Thus, increasing efforts are being made in assessing the usage of pedestrian infrastructure, identifying critical features such as pedestrian density (here referring to the number of pedestrians per unit area of sidewalk space) and pedestrian speed, to better understand the complex pedestrian mobility patterns and interactions (Fruin,, 1970; Maghelal and Capp,, 2011; Dragović et al.,, 2023). Currently, pedestrian flow data is typically collected through infrared (IR) break-beam sensors and cameras placed at specific locations across the city (Moussaïd et al.,, 2010; Alahi et al.,, 2014; Carter et al.,, 2020). While this type of solution can offer granular data, it is typically difficult to adopt it on a broad scale due to its high costs and privacy issues (e.g., General Data Protection Regulation in Europe). More recently, street-level imagery (e.g., Google Street View) has been explored as an alternative source, where pedestrians can be detected and counted with computer vision techniques (Liu et al.,, 2023; Chen et al.,, 2020). These imagery-based approaches are less costly to deploy and scalable across large areas, but they suffer from limited temporal resolution, since imagery is updated infrequently and cannot provide continuous measurements. Other studies focusing on the larger-scale analysis of urban mobility patterns (e.g., origins and destinations, mode choice, trip features) rely on GPS, wireless and cellular networks, and social media geotag records (Leslie et al.,, 2007; Kung et al.,, 2014; D’Orso and Migliore,, 2020; Gao et al.,, 2014). These solutions offer a good overview of pedestrian mobility at an aggregate level, but they cannot capture more complex phenomena such as, pedestrian flows, trajectories, and interactions with other road users. Surveys and audits are also used for collecting qualitative data on perceived walkability (Steve and Shane,, 2011; Cambra and Moura,, 2020; Chen et al.,, 2024) and quantitative measures of the built environment to develop walkability indicators (Bartzokas-Tsiompras et al.,, 2021; Lee et al.,, 2020). Although these methods provide useful insights, they are time-consuming, resource-intensive, and difficult to scale, posing a major bottleneck to advancing the knowledge of walkability in research.
@@
This study does not aim to assess all variables related to walkability. Instead, it demonstrates how sidewalk robots can be leveraged to automatically capture measurable walkability-related features at the sidewalk level and use them to assess pedestrian mobility in practice. The contributions of this study are threefold. First, it introduces robots as a novel data collection instrument, capable of overcoming limitations of traditional methods which prevent their application and scalability. As sidewalk delivery robots become more common in urban environments, their routine movements offer a practical and cost-efficient means to evaluate pedestrian mobility real-time. Second, by grouping different operational, walkability-related factors into robot trip features, sidewalk condition features, and sidewalk utilization features, the study establishes a systematic framework for extracting and analyzing data. Third, it provides empirical evidence that robot-based measurements can both confirm established relationships in pedestrian dynamics and reveal context-specific patterns, such as how width, density, and surface quality shape walking speed and movement styles. In the long run, this robot-based method can help build richer and more inclusive datasets that complement traditional audits and fixed sensors.
@@ 2 Related Work
While these studies offer a broad conceptual and empirical understanding of walkability, this study focuses on the operational features, specific environmental factors that directly affect the functional experience of walking. Rather than a systematic survey of the entire walkability literature, we summarize relevant across complementary fields, including urban planning, transport geography, pedestrian mobility studies, and traffic flow theory. The selected studies explicitly examined the relation between the built environment and pedestrian movements, provided measurable indicators at the sidewalk level, and offered empirical insights. On this basis , we identified twelve relevant factors from an operational perspective: 1) static obstacles; 2) dynamic obstacles; 3) walkway width; 4) walkway quality; 5) sidewalk slope; 6) crosswalk distance; 7) crosswalk type; 8) potential for conflicts; 9) vehicle speed; 10) traffic volume; 11) driveways; 12) supporting facilities. Among these, factors 6 to 12 are mainly related to crossing conditions. Each feature may correspond to multiple indicators or measurements across different studies.
@@ 3.1 Experimental setup
From September 18, 2024, to March 24, 2025, the robot completed 101 trips, covering 900 records of sidewalk segments in total. Trips were conducted on different days and at various times between 08:00 and 18:00, with particular focus on peak hours. During each trip, onboard sensors continuously recorded raw data, including GNSS points, robot velocity and heading, IMU readings, sidewalk width, ambient brightness, and anonymized pedestrian data. The experiments were conducting during months of higher pedestrian activity on campus, ensuring the dataset’s representativeness.
The recorded raw data from each trip were stored in a centralized data warehouse that maintains all historical trip records. Each trip was divided into sidewalk segments, and segment-level features were extracted and grouped into three categories: robot trip features, sidewalk condition features, and sidewalk utilization features. These extracted features offer valuable insights into the quality of walking infrastructure and spatio-temporal pedestrian patterns. An overview of these features is presented in Table 2, while an interactive dashboard [^1] https://ismir-sidewalk-mobility-data-dashboard.streamlit.app/ visualizes sidewalk usage and characteristics for the study area. In the subsequent analyses (Section 4), these features are used in two ways: (i) segment-wise averages are computed to compare conditions and utilization across the network; and (ii) the individual segment-level records are analyzed directly for box-plot visualization, Pearson correlation, clustering, and regression.
@@ 3.4 Sidewalk utilization features
4 Results
We perform a series of analyses on the extracted sidewalk features to investigate their walkability-related implications. All analyses focus on sidewalks and exclude pedestrian crossings, as crossings primarily reflect intersection control, signal timing, and traffic exposure rather than sidewalk conditions or pedestrian mobility along continuous paths. Subsection 4.1 presents an exploratory analysis of sidewalk condition features, highlighting variation in surface quality, width, and slope across different segments. Subsection 4.2 shifts the focus to aggregated pedestrian utilization patterns like speed and density relations. Building on these analyses, Subsection 4.3.1 applies clustering to identify distinct pedestrian behavior categories and analyze their relation to environmental and infrastructure characteristics. Finally, Subsection 4.3.2 employs regression to investigate the potential influence of sidewalk operational features on pedestrian movement.
@@ 4.1 Sidewalk Condition Variability Across the Network
In contrast, the irregularity index exhibits higher variability across the network, with Segment 7 displaying a significant spike due to frequent paving pattern changes (see Figure 7(b)). When the robot travels through this segment, it encounters four abrupt paving transitions within a short distance, including a cycling lane in between. Segment 7 records the highest irregularity but only moderate unevenness, indicating that the two IMU-based metrics capture complementary aspects. Irregularity is also elevated on Segment 8 (manhole cover, storm drain), whereas Segments 2–3 are minimal. Together, these results confirm that our proposed unevenness and irregularity metrics accurately describe actual sidewalk conditions, providing valuable insights for assessing pedestrian comfort and sidewalk robot mobility.
The distributions of detected average effective walkway width during the robot’s trips across various sidewalks are shown in Figure 7(c). Narrow sidewalks (Segments 1-4 and 7) display tight width distributions, generally below 3 m, while open areas (around 5.5–10 m) lack clear edges and show greater variability. This variation reflects both geometric differences and the robot’s sensing limitation. The stereo camera which offers a 120-degree field of view, in practice captures sidewalk information up to a maximum of about 7 meters ahead when centered. In broader or less structured spaces, the robot’s steering to avoid pedestrians or obstacles can further underestimate or fluctuate measured width. Consequently, width estimates are most reliable in narrow, well-defined corridors, while some variation in these areas may still result from temporary obstacles encountered during passage.
Figure 7(d) presents the absolute slope values for each sidewalk segment, capturing steepness differences relevant to walkability and robot navigation. Segment 7 is the steepest, whereas Segments 1 and 2 are nearly flat, reflecting modest but notable variations across the network.
Overall, surface quality, effective width, and slope show clear segment-level variability that shapes the walking environment and provides the basis for the utilization and behavioral analyses in subsequent sections.
@@ 4.2 Sidewalk Utilization and Density-Speed Dynamics
Average pedestrian speed and pedestrian density are estimated during each robot traversal of the sidewalk segments. The distributions of these two features across different sidewalks are presented in Figure 8(a) and Figure 8(b). The results reveal notable variability in pedestrian behavior across segments. Segments 3 and 4 show broader speed distributions, suggesting a mix of movement patterns and variable flow conditions, while Segments 6 and 9 display more concentrated speeds, indicating steadier movement.
Pedestrian density also differs across sidewalks. Although Segments 6 and 7 are located on opposite sides of the same road, Segment 6 shows the highest density, while Segment 7 remains sparse. As shown in Figure 7(c), their average sidewalk widths are approximately 6 m and 2 m, respectively, illustrating how available space influences pedestrian volumes. Segment 6 serves as the main access corridor connecting the metro and light rail station exits to the campus entrance, and it is lined with convenience stores and restaurants. By contrast, segment 7 lacks notable points of interest, resulting in limited pedestrian activity.
@@
The observed densities in our dataset are relatively low ( up to 0.34 peds/m2), while walking speeds reach 2.5 m/s, reflecting the open campus setting where pedestrians move purposefully in an outdoor environment. Despite these scale differences, the overall pattern follows the expected inverse relationship between density and speed: higher densities correspond to slower movement. Measurement errors due to the robot’s onboard hardware and perception-limited capabilities may also contribute to such discrepancy. The robot relies on real-time pedestrian-detection algorithms (Jocher et al.,, 2023), which are constrained by onboard computational power and may not achieve the precision of offline or stationary sensing systems.
At low densities (below 0.1 peds/m2), speeds vary widely from under 0.5 m/s to over 2 m/s, reflecting heterogeneous walking behaviors. As density increases beyond 0.15 peds/m2, speeds converge below 1.5 m/s, suggesting the onset of congestion and greater interaction among pedestrians. Segment-level variations also emerge. For instance, Segment 4 exhibits a broader spread of speeds, while Segment 1 consistently supports higher walking speeds even at moderate densities. These patterns suggest that both spatial context and sidewalk characteristics influence how pedestrians adapt their pace in shared environments.
The analysis confirms the expected inverse relationship between pedestrian density and speed, while also revealing context-specific variations across sidewalks. Together, these findings characterize how pedestrians collectively respond to spatial constraints and flow conditions, offering empirical grounding for the behavioral analyses that follow.